a disc of mass 5kg is kept floating horizontally in mid air by firing 10 bullets per seconds vertically up. If the mass of the bullets is 50 grams and the bullets rebound with the same speed,the speed of each bullet is

Grade:11

1 Answers

Aman
7 years ago
 
let speed be v.
the forces acting on the disc are its weight(downwards)and the force due to change in momentum of the bullet transferred to it.
Change in momentum=mv-(-mv)=2mv
Force = (change in momentum)(time)=2mvt=2(50/1000)(v)(1)=v/10
As the disc is in vertical equilibrium,
mg=v/10
(5)(10)=v/10
v=500 m/s